Assume that a randomly selected subject is given a bone density test. Those test scores are normally distributed with a mean of 0 and a standard deviation of 1. Find the probability that a given score is less than −1.56 and draw a sketch of the region.
Answer #1

Any Let X denotes test scores XNN 14-0,62=1) then P(x2-1:56) = ? PIX-02 -1:560) = P(Z 2-1-56) 1 =171-1.56) PIX_21156) = (-156